x+3y=1 -4y=x find x
Answer by jim_thompson5910(35256)

-4y = x is the same as x = -4y

Start with x+3y=1 and replace every x with -4y. Then solve for y

x+3y=1

-4y+3y=1

-y = 1

y = -1

Then use this to find x


x = -4y

x = -4(-1)

x = 4


So x = 4 and y = -1


------------------------------------------------------------------------

Let me know if that helps or not. Thanks.

If you need more help, feel free to email me at jim_thompson5910@hotmail.com

My Website: http://www.freewebs.com/jimthompson5910/home.html